You are standing in the middle of a large room listening to a cacaphony of sounds. Rank the intensity level of the sounds from each source (1-loudest, 2-next loudest, ...) Consider all objects to be right next to you. If two values are wthin 0.1 dB from each other, consider them to be the same. Ten buzzers each producing a sound with intensity level of SOdB Four buzzers each producing a sound with intensity level of 54dB. One hundred buzzers each producing a sound with intensity level of 40dB An object emitting a sound with an intensity level of 60dB. Two buzzers each producing a sound with intensity level of 57dB. Submit Answer Tries 0/10
Answer #1

1) Ten buzzers each producing a sound with intensity level of 50 dB.
10 log (10 x (10^5)) = 60 dB

2) 4 buzzers each producing a sound with intensity level of 54 dB.
10 log (4 x (10^5.4)) = 60.02 dB

3) 100 buzzers each producing a sound with intensity level of 40 dB.
10 log (100 x (10^4)) = 60 dB

4) An object emitting a sound with an intensity level of 60 dB


5) 2 buzzers each producing a sound with intensity level of 57 dB.
10 log (2 x (10^5.7)) = 60.01 dB

The order will be

2) > 5) > 1) = 3) = 4)

Other way, all sounds are almost equal.